Find the maximum value of force F that can be applied to the system of the rod AB (Area=2 in?, E=29x106 psi) and the rigidEMB=0 hinge B 5 4 6
Answer #1

Lemgth of AB=12"

Area of AB=2 in²

E of AB=29000 ksi

Max stress in AB=20 Ksi

Max load allowable in AB=20*2=40 kips

Verticak displacemet of B=(20/29000)*12=0.00827 in

Vertical displacement at D=0.00827*(6+4)/6=0.0137 in<0.05 in

Therefore, F should be such that force in AB is 40 kips

From moment equilibrium about C,

F*4=40*6

F=60 kips

This is the required answer